The density of states of electron (including spin degeneracy) in the band is given by
1
Expert's answer
2021-10-05T10:08:25-0400

The density of states of electron is given by


"g(k)\text dk=\frac{k'^2}{\pi^2}\sqrt{\frac{m_xm_ym_z}{m'^3}}\text dk'."